OpenStudy (anonymous):

evaluate log of 9 71

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (immanuelv):

1.940029217

OpenStudy (immanuelv):

is it log base 9?

OpenStudy (anonymous):

\[log_9(71)=\frac{ln(91)}{ln(9)}\]

OpenStudy (immanuelv):

uhhh. u mean ln 71?

OpenStudy (anonymous):

i meant \[\frac{ln(71)}{ln(9)}\]

OpenStudy (anonymous):

silly me

OpenStudy (anonymous):

are you sure it isn't \[log_9(81)=2\] because \[9^2=81\]?

OpenStudy (anonymous):

you guys rock!

OpenStudy (anonymous):

yes its \[\log_{9} 71\]

OpenStudy (anonymous):

k then use change of base formula as we wrote

OpenStudy (anonymous):

change to what

OpenStudy (anonymous):

btw you can use \[\frac{log(71)}{log(9)}\] too, it doesn't matter what base you use

OpenStudy (anonymous):

oh okayy i get what your saying so the first one is right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

91 was a typo

OpenStudy (anonymous):

yeah